A 14kg is being pulled with a force of 40N at a 22o angle. What is the acceleration if coefficient of friction is 0.35?


1
Expert's answer
2021-04-14T20:59:52-0400

By defintion, the friction force is:


"F_{fr} = \mu mg\\
F_{fr} = 0.35\cdot 15\cdot 9.8 = 51.45N"

where "\mu = 0.35" is the coefficient of friction, "m = 14kg" is the mass of the body, and "g = 9.8 m\/s^2" is the gravitational acceleration.

The horizontal projection of the pulling force is:


"F_x = F\cos \theta\\
F_x = 40N\cdot \cos 22\degree \approx 37N"

where "F = 40N", and "\theta = 22\degree".

Since the friction force is greater than the pulling force:

"F_{fr} > F_x"

the body will not move (the acceleration will be 0).


Answer. 0.